OpenStudy (anonymous):

vg=-443j+56.56i find the actual speed

OpenStudy (perl):

we need more information to solve this

OpenStudy (anonymous):

Unless there's any more information, I'm guessing the way you'd solve this is to find the resulting vector using the formula:\[r^2=x^2+y^2 = (-443)^2+56.56^2=196249+3203.56=199452.56\]And hence \[r=\sqrt{199452.56}=446.601\]

OpenStudy (anonymous):

Perl you assisted me with the first part of the question, here is the second part Find the velocity vg of the 737 relative to the ground. (Hint: vg = va + vw) i did this to find the velocity, Vg=Va+Vw Vg=(-500j+56.56j)+56.56i and i got Vg=-443j+56.56i with this we need to find the actual speed

OpenStudy (perl):

ok one sec

OpenStudy (anonymous):

my instructor gave us a hint \[\left| \left| Vg \right| \right|\]

OpenStudy (perl):

so we had <0, -500> + <56.67 , 56.67> = < 56.67 , -443.33>

OpenStudy (anonymous):

would toms answer be correct

OpenStudy (perl):

vg = va + vw = <0, -500> + <56.5 , 56.57> = < 56.57 , -443.33> | vg | = sqrt ( 56.57 ^2 + 443.33^2)

OpenStudy (perl):

and the 'actual speed' is the ground speed

OpenStudy (anonymous):

i have a few more if ur interested in helping with